OpenStudy (anonymous):

What is the equation, in standard form, of a parabola that models the values in the table? x -2, 0, 4 f(x) 2, -2, 86

OpenStudy (amistre64):

solve the system of equations: (-2)^2a -2b +c = 2 (0)^2a +0b +c = -2 (4)^2a +4b +c = 86

OpenStudy (amistre64):

well, since the middle gives us c=-2, its actually 2 equations in 2 unknowns left to determine

OpenStudy (anonymous):

I am completely lost. Is there a way you can explain to me how to do this problem?

OpenStudy (amistre64):

we should know that a general quadratic is of the form: f(x) = ax^2 + bx + c in order for those 3 points to satisfy some abc construction of this form, we simply insert the x,y values (-2,2): 2 = a(-2)^2 + b(-2) + c (0, -2):-2 = a(0)^2 + b(0) + c (4, 86): 86 = a(4)^2 + b(4) + c --------------------------------------- this forms 3 equations in 3 unknowns. 2 = 4a - 2b + c -2 = 0a + 0b + c 86 = 16a + 4b + c ------------------------------------ it is obvious from the middle equation that c=-2, this reduces the setup to 2 equations in 2 unknowns 2 = 4a - 2b - 2 86 = 16a + 4b - 2 or written another way: 4 = 4a - 2b 88 = 16a + 4b continue solving for a and b